OpenStudy (anonymous):

What is √176 factored. I know the roots come down to 2*2*2*11 but I don't know what to do from there.

OpenStudy (anonymous):

You are very close and to be commended for trying and for getting so close. Actually, when you factor 176, you get: 2 x 2 x 2 x 2 x 11 So, sqrt(2 x 2 x 2 x 2 x 11) = sqrt(4 x 4 x 11) = sqrt(4^2 x 11) = sqrt(4^2) x sqrt(11) = 4 x sqrt(11)

OpenStudy (anonymous):

Thx for the recognition! More importantly, good luck to you in all of your studies!

OpenStudy (anonymous):

@tcarroll010 thanks so much!

OpenStudy (anonymous):

You are quite welcome! Have a nice day!